• Simplifying the expression 5/6-3/4+7/2 gives?
1
Expert's answer
2021-04-28T07:24:48-0400
"\dfrac{5}{6}-\dfrac{3}{4}+\dfrac{7}{2}=\dfrac{10-9+42}{12}=\dfrac{43}{12}"